Given an integer

N as input, write a program to print a number diamond of 2*N -1 rows as shown below.Note: There is a space after each number.

Input

The first line of input is an integer

N.

In the given example, the number of rows in the diamond is

5.So, the output should be

    1 
   1 2 
  1 2 3 
 1 2 3 4 
1 2 3 4 5 
 1 2 3 4 
  1 2 3 
   1 2 
    1

4,So the output should be

   1 
  1 2 
 1 2 3 
1 2 3 4 
 1 2 3 
  1 2 
   1
1
Expert's answer
2021-07-27T07:43:18-0400
rows = int(input("Enter the number of rows: "))  
  
k = 2 * rows - 2  
for i in range(0, rows):  
    for j in range(0, k):  
        print(end=" ")  
    k = k - 1  
    for j in range(1, i + 1):  
        print(j, end=" ")  
    print("")  
  


k = rows - 2  
for i in range(rows, -1, -1):  
    for j in range(k, 0, -1):  
        print(end=" ")  
    k = k + 1  
    for j in range(1, i + 1):  
        print(j, end=" ")  
    print("")